OpenStudy (anonymous):

how do i convert a decimal (integer) to a hexadecimal? To an octal? to a binary?

OpenStudy (anonymous):

unsigned integer

OpenStudy (anonymous):

best explains by example 42 decimasl = 42 / 8 = 5 renainder 2 = 52 octal 42 decimal = 42/16 = 2A hexadecimal the digits i hexadecimal range from 0 -9 then A,B,C,D E F

OpenStudy (anonymous):

how do I convert 328789 to binary?

OpenStudy (radar):

42 decimal to binary 32 16 8 4 2 1 1 0 1 0 1 0

OpenStudy (anonymous):

in binary there is only 2 digits 0 and 1 so 10 = 2^1 + 0 = 2 decimal 11 = 2^1 + 1 = 3 decimal 111 = 2^2 + 2^1 + 2^0 = 4 + 2 +1 = 7 decimal

OpenStudy (anonymous):

to get 328789 - extend radar's series to left until you have values high enough.

OpenStudy (anonymous):

328789 is a long number. 0b1010000010001010101

OpenStudy (anonymous):

it's only 19 bits... am I right?

OpenStudy (anonymous):

yes - i just wolframalpha lol - eaiest way

OpenStudy (anonymous):

I should learn how to do it by hand/head, since I will be using bits and octets a lot.

OpenStudy (anonymous):

yea

OpenStudy (radar):

A hint if the number is odd the LSB is a 1 lol

OpenStudy (anonymous):

yes.